FOURIER TRANSFORM EMISSION SPECTROSCOPY OF A NEW ^{3}Φ − a^{3}Δ SYSTEM OF ScH AND ScD

\underline{R.S. RAM}, Department of Chemistry, University of Arizona, Tucson, AZ 85721; P. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada N2L 3G1.

The emission spectra of ScH and ScD have been investigated at high resolution in the 3000-14500 cm^{−1} region. The molecules were excited in a scandium hollow cathode lamp by discharging a mixture of He and H_{2} or D_{2}, and the spectra were recorded using a Fourier transform spectrometer. The new bands of ScH and ScD with high wavenumber R heads near 11620 cm^{−1} and 11630 cm^{−1}, respectively, have been assigned as the 0-0 bands of the ^{3}Φ − a^{3}Δ transition. The rotational structure of each sub-band consists of only the R and Q branches. A rotational analysis of these bands have been obtained and the spectroscopic constants have been extracted. This transition is analogous to the^{3}Φ−a^{3}Δ transition of YH [Ram and Bernath, J. Chem. Phys., 101, 9283 (1994)] and LaH [Ram and Bernath, J. Chem. Phys., 104, 6444 (1996)].